OpenStudy (anonymous):

assume the y varies inversely with x and y=-3 when x=4. Write an indirect variation equation for the relationship.

OpenStudy (anonymous):

If it varies inversely that means y=k/x. Now plug in your values -3=k/4. So k=-12. That means y=-12/x